OpenStudy (anonymous):

generalize the pattern by finding the nth term beggining with n=1 16,8,4,2

OpenStudy (cwrw238):

you have a geometric series here if first term =a and common ratio= r, the nth term = a r^(n-1)

OpenStudy (cwrw238):

her the common ration = second term / first term = 8 / 16 = 0.5

OpenStudy (cwrw238):

note of course that 4/8 and 2/4 = 0.5 also

OpenStudy (cwrw238):

each term is half of the previous one

OpenStudy (cwrw238):

so next 3 terms are 1, 0.5 ,0.25

OpenStudy (cwrw238):

if you wanted to find say the 50th term you could keep on dividing by 2 bur it would be easier to use the formula and your calculator: 50th term = 16* (0.5)^(50-1) = 2.842170943 * 10^-14

OpenStudy (cwrw238):

so the formula is nth term = a r (n-1) where a = 16 and r = 0.5 that is; nth term = 16 * (0.5)^(n-1)